# Fast post-process Bayesian inference with Variational Sparse Bayesian Quadrature

### v3

@@ 3.3 Noise shaping
Figure 3: Illustration of noise shaping effect. Noise shaping improves the selection of inducing points and the approximation of a sparse GP in the high posterior density region. A better sparse GP consequently leads to an improved variational posterior. The red triangles are locations of inducing points selected via Eq. 25. The orange line represents the sparse GP prediction. The green dashed line represents the variational posterior. (a) True density function (black line) and the observations (black circles). (b) Sparse GP approximation to the log-density function without noise shaping. The black curve is the ground truth log-density function. (c) Variational posterior without noise shaping. (d) Noise shaping curve corresponding to Eq. 3.3. (e) Sparse GP approximation with noise shaping. (f) Variational posterior with noise shaping.
@@
where $q(\mathbf{f})$ is the variational GP posterior. Thus, with noise shaping the expected log-likelihood term of the GP-elbo becomes,
@@
where $w_{n}\equiv\frac{\sigma^{2}_{\text{obs}}(\mathbf{x}_{n})}{\sigma^{2}_{\text{tot}}(\mathbf{x}_{n},y_{n})}\leq 1$. That is, by assigning larger ‘shaping’ noise to low-density observations we are downweighing their role in the sparse GP representation, guiding the sparse GP to better represent higher-density regions. Noise shaping also helps during inducing point selection (Eq. 25), as shown in Figure 3.
@@ 3.4 Approximation error
We introduce Lemma 3.1 and 3.2 first, whose proofs are provided in Supplementary Material B. Recall that $\mathbf{f}$ denotes the latent function values corresponding to the observations $\mathbf{y}$, $\mathbf{u}$ represents values at the inducing points $\mathbf{Z}$, and $\bm{\psi}$ denotes the optimal variational parameters of the sparse GP. Denote with $f$ and $f_{e}$ the posterior predictive functions of the sparse GP and exact GP, respectively. The posterior mean functions of the sparse GP and exact GP are represented as $\bar{f}$ and $\bar{f}_{e}$. Finally, $\|\cdot\|_{\text{TV}}$ denotes the total variation distance.
@@
Assume that $D_{\text{KL}}(p(\mathbf{f},\mathbf{u}\mid\mathbf{y})\|p(\mathbf{f},\mathbf{u}\mid\bm{\psi}))<\gamma$. Then, for any $\ell>0$ there exists $K_{\ell}$ such that, for any $\mathbf{x}^{*}$, $\|\mathbb{E}[{f}(\mathbf{x}^{*})^{\ell}]-\mathbb{E}[{f}_{e}(\mathbf{x}^{*})^{\ell}]\|<K_{\ell}\sqrt{\gamma/2}$. There also exists $K_{\mathrm{e}}$ such that, for any $\mathbf{x}^{*}$, $\|\mathbb{E}[\mathrm{exp}({f}(\mathbf{x}^{*}))]-\mathbb{E}[\mathrm{exp}({f}_{e}(\mathbf{x}^{*}))]\|<K_{\mathrm{e}}\sqrt{\gamma/2}$.
@@ 4.1 Baseline methods
We recall that with black-box inference we mean that the target lacks gradients and may be expensive to evaluate and possibly noisy. As few other methods afford post-process and black-box inference of the posterior, we compare vsbq again three baselines: black-box variational inference (bbvi), neural network regression (nnr) and the popular Laplace approximation (laplace). Of these, only nnr is also a post-process method. Further implementation details of vsbq are provided in Supplementary Material C.1.
Black-box variational inference.
When the target posterior density is a black-box, the reparametrization trick (Kingma and Welling, 2013) cannot be applied to estimate the elbo gradient for stochastic variational inference. Instead, one needs to resort to other techniques for computing the gradient of the elbo, such as the score function estimator, also known as the REINFORCE estimator (Ranganath et al., 2014), which often has higher variance compared to the reparametrization trick (Gal, 2016; Xu et al., 2019). Specifically, by differentiating Eq. 2, the elbo gradient $\nabla_{\bm{\phi}}\textsc{elbo}\left[q_{\bm{\phi}}\right]$ can be written as,
@@
Neural network regression.
For a direct comparison with our method, we develop a post-process inference algorithm based on a deep neural network regression surrogate (nnr) instead of a sparse GP surrogate, otherwise leaving the post-process procedure (Figure 2) as much as possible the same. This is a competitive baseline since deep neural networks exhibit strong regression performance in the presence of a large number of data points (Goodfellow et al., 2016). For the network architecture, we use a multilayer perceptron (MLP) with an input layer of dimension $D$, four hidden layers of $1024$ units, and an output layer for predicting the log-density value. The activation function is chosen to be the rectified linear units (ReLu; Goodfellow et al., 2016, Chapter 6). In addition, we add a negative quadratic mean function to the neural network output to ensure that it represents a valid log-density surrogate function. [^7] An MLP with ReLu activations is a continuous piecewise affine function (Mukherjee et al., 2018), and therefore adding a trainable negative quadratic mean function ensures the integrability of the exponentiated surrogate. The negative quadratic mean function is the same as the one used for the (sparse) GP (see Eq. 17). In total, the surrogate function $g$ is:
@@
$\mathcal{L}(\mathbf{w})=\sum_{n=1}^{N}\frac{\left(g(\mathbf{x}_{n})-y_{n}\right)^{2}}{\sigma_{\text{tot}}^{2}(\mathbf{x}_{n},y_{n})},$ (35)
where $\sigma_{\text{tot}}^{2}(\mathbf{x}_{n},y_{n})=\sigma_{\text{obs}}^{2}(\mathbf{x}_{n})+\sigma_{\text{shape}}^{2}(\Delta y_{n})$, as per noise shaping. [^8] We found that noise shaping empirically helped stabilize the neural network training. We optimize the neural network parameters by minimizing the objective in Eq. 35 with the AdamW optimizer (Loshchilov and Hutter, 2019). For regularization, we considered the ‘weight decay’ hyperparameter $\alpha\in\{0,0.01,0.1\}$, selecting the best neural network surrogate based on the loss on a split validation dataset. Finally, we use stochastic variational inference with the reparametrization trick (Kingma and Welling, 2013) to compute the approximate posterior. This part is the same as vsbq except that the expected log joint of the surrogate in Eq. 2 is approximated via Monte Carlo samples rather than calculated exactly via sparse Bayesian quadrature. For more details, see Supplementary Material C.4.
Laplace approximation.
@@ 4.2 Procedure and metrics
Procedure.
The optimization trace points and corresponding log-density values are collected as the training dataset for vsbq and nnr. For each problem, we repeated the entire optimization procedure – each with multiple MAP estimation runs, as explained above – ten times with different random seeds. This yielded ten different training sets per problem, used to assess the robustness and reliability of the methods. The number of inducing points for the sparse GP is set to $100D$. For bbvi, as mentioned in Section 4.1, a budget of $10\times 3000D=30000D$ target density evaluations per random seed is allocated for stochastic optimization. In the case of a noisy target, we further vary the observation noise level to study the noise sensitivity of the methods. For the Laplace approximation, we first find the MAP point by transforming the parameter space to unbounded if needed, via a nonlinear mapping. [^9] For the Laplace approximation to be valid, it is necessary to find the MAP estimate in the unbounded space subsequently used to compute the multivariate normal approximation; this is particularly important if the mode is close to the bounds. We subsequently compute the Hessian matrix via adaptive numerical differentiation (Brodtkorb and D’Errico, 2022). Finally, we compute the performance metrics as detailed below.
Metrics.
@@
    where $D_{\text{KL}}\left(p||q\right)$ is the Kullback-Leibler divergence between distributions $p$ and $q$ and $\mathcal{N}[p]$ is a multivariate normal distribution with mean equal to the mean of $p$ and covariance matrix equal to the covariance of $p$ (and same for $q$). Eq. 37 can be expressed in closed form in terms of the means and covariance matrices of $p$ and $q$. For reference, two Gaussians with unit variance and whose means differ by $\sqrt{2}$ (resp., $\frac{1}{2}$) have a GsKL of 1 (resp., $\frac{1}{8}$). As a rule of thumb, we consider a desirable target to have GsKL less than $\frac{1}{8}$.
@@ 4.3 Synthetic problems
Two Moons bimodal posterior.
@@
Multivariate Rosenbrock-Gaussian.
@@ 4.4 Real-world models
Noisy likelihood evaluations.
@@
Bayesian timing model.
@@
Figure 5: Bayesian timing model. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). vsbq performs well across all noise levels and can outperform (noiseless) laplace even under high log-likelihood noise, whereas nnr demonstrates less robustness with several failed runs. bbvi methods exhibit similar performance above the metric thresholds.
Multisensory causal inference.
@@
Figure 6: Multisensory causal inference model. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). vsbq performs well across low to moderate noise levels and outperforms (noiseless) laplace even at high noise levels. nnr also demonstrates strong performance across all noise levels. bbvi methods perform similarly and are above the thresholds, except for the MoG($K=50$) at low noise levels.
@@ 4.5 Additional analyses
We summarize here the results of a number of additional analyses reported in the Supplementary Material.
MAP estimation with Bayesian Adaptive Direct Search (BADS).
@@
In the earlier paragraphs, we presented results obtained by approximating the target posterior density from the surrogate of the log density (Figure 1) via variational inference. A natural alternative is to instead run mcmc on the surrogate log-density (Rasmussen, 2003; Nemeth and Sherlock, 2018; Järvenpää et al., 2021; Gammal et al., 2022). In Supplementary Material C.8, we provide experimental results showing that mcmc can underperform in this setting and discuss the reasons.
@@ 5.2 Limitations and future work
Arguably, post-process inference via vsbq is constrained to low-dimensional problems for several reasons: as the dimension increases, pre-existing evaluations from MAP optimizations become less likely to cover the majority of the posterior mass (Vershynin, 2018, Chapter 3.3.3). Moreover, the MAP estimate may become not meaningful or even ill-defined in certain settings, e.g., when the likelihood is unbounded (Gelman et al., 2013). While a larger number of evaluations might expand the range of tractable dimensions, the curse of dimensionality poses challenges in approximating a function without additional structural assumptions. A potential solution to address the issue of lack of coverage is active sampling, that is the ability to acquire new log-density evaluations where needed (Osborne et al., 2012; Acerbi, 2020; Järvenpää et al., 2021; De Souza et al., 2022). Active sampling would need one or multiple rounds of ad-hoc function evaluations and subsequent surrogate model updates, which would substantially increase the post-processing time (De Souza et al., 2022). Finding efficient methods for active learning for post-process inference is left for future research. Despite the above limitations, the relatively cheap cost of our method makes it still valuable for quickly constructing a tractable initial approximation of the posterior, potentially useful to inform subsequent runs of mcmc or other inference methods (Zhang et al., 2022).
@@
An important feature of approximate inference methods is the availability of diagnostics to assess the reliability of inference (Vehtari et al., 2021; Yao et al., 2018). A general-purpose inference diagnostic consists of posterior-predictive checks, i.e., testing that synthetic data generated from parameters sampled from the posterior are compatible with the actual data (Gelman et al., 2013). As an additional diagnostic, vsbq provides the standard deviation of the elbo, $\textsc{elbo}_{\text{sd}}$, which can be calculated via sparse Bayesian quadrature (Eq. 23). $\textsc{elbo}_{\text{sd}}$ reflects the uncertainty of the sparse GP prediction in regions where the variational posterior has non-negligible mass. Therefore, it can serve as a useful diagnostic in that solutions with $\textsc{elbo}_{\text{sd}}\gg 1$ should not be trusted. However, as is often the case with inference diagnostics, a small $\textsc{elbo}_{\text{sd}}$ does not guarantee the validity of the approximate variational posterior (Acerbi, 2018). As a practical validation approach, we recommend running vsbq multiple times by dropping for example 20% of the training set and checking the consistency of the approximate posteriors via a form of cross-validation. Visualization of the posterior together with the locations of the evaluated points is also helpful for validating that the approximate posterior is supported by actual evaluations, as opposed to escaping from the training points region, leading to ‘hallucinated’ posterior regions (De Souza et al., 2022). We further discuss this ‘hallucination’ problem in Supplementary Material C.8. Finally, if additional exact log-density evaluations are possible, one can leverage Pareto smoothed importance sampling (Vehtari et al., 2015) for correcting and validating the approximate posterior (Yao et al., 2018).

### v4

@@ 3.3 Noise shaping
(a) True density function (black line) and the observations (gray circles).
(b) Sparse GP approximation (orange line) to the true log-density function (black line) without noise shaping.
(c) Variational posterior (green dashed line) without noise shaping.
(d) Noise shaping curve corresponding to Eq. 3.3.
(e) Sparse GP approximation with noise shaping.
(f) Variational posterior with noise shaping.
Figure 3: Illustration of noise shaping effect. The red triangles are locations of inducing points selected via Eq. 25. Noise shaping (bottom row) improves the selection of inducing points and the approximation of the sparse GP in the high posterior density region, compared to a sparse GP without noise shaping (top row). A better sparse GP consequently leads to an improved variational posterior.
@@
where $q(\mathbf{f})$ is the variational GP posterior. Thus, with noise shaping, the expected log-likelihood term of the GP-elbo becomes,
@@
where $w_{n}\equiv\frac{\sigma^{2}_{\text{obs}}(\mathbf{x}_{n})}{\sigma^{2}_{\text{tot}}(\mathbf{x}_{n},y_{n})}\leq 1$. That is, by assigning larger ‘shaping’ noise to low-density observations, we are downweighing their role in the sparse GP representation, guiding the sparse GP to better represent higher-density regions. Noise shaping also helps during inducing point selection (Eq. 25), as shown in Figure 3.
@@ 3.4 Approximation error
We introduce Lemma 3.1 and 3.2 first, whose proofs are provided in Supplementary Material B. Recall from Section 2.6 that $\mathbf{f}$ denotes the latent function values corresponding to the observations $\mathbf{y}$, $\mathbf{u}$ represents values at the inducing points $\mathbf{Z}$, $\bm{\psi}$ denotes the optimal variational parameters of the sparse GP and $p(\mathbf{f},\mathbf{u}\mid\bm{\psi})$ is the joint posterior distribution of the sparse GP given the optimal variational parameters. Denote with $f$ and $f_{e}$ the posterior predictive functions of the sparse GP and exact GP, respectively. The posterior mean functions of the sparse GP and exact GP are represented as $\bar{f}$ and $\bar{f}_{e}$. Finally, $\|\cdot\|_{\text{TV}}$ denotes the total variation distance.
@@
Assume that $D_{\text{KL}}(p(\mathbf{f},\mathbf{u}\mid\bm{\psi})\|p(\mathbf{f},\mathbf{u}\mid\mathbf{y}))<\gamma$. Then, for any $\ell>0$ there exists $K_{\ell}$ such that, for any $\mathbf{x}^{*}$, $|\mathbb{E}[{f}(\mathbf{x}^{*})^{\ell}]-\mathbb{E}[{f}_{e}(\mathbf{x}^{*})^{\ell}]|<K_{\ell}\sqrt{\gamma/2}$. There also exists $K_{\mathrm{e}}$ such that, for any $\mathbf{x}^{*}$, $|\mathbb{E}[\mathrm{exp}({f}(\mathbf{x}^{*}))]-\mathbb{E}[\mathrm{exp}({f}_{e}(\mathbf{x}^{*}))]|<K_{\mathrm{e}}\sqrt{\gamma/2}$.
@@
Theorem 3.3 provides a theoretical justification for the quality of the variational posterior obtained from a sparse GP surrogate. While the bound is not directly useful for empirical tuning—due to its dependence on constants such as $K_{l}$ and $\gamma$—it reveals how the total error decomposes into two main interpretable and controllable sources: the error introduced by approximating the exact GP with a sparse GP, and the error due to variational inference over the surrogate.
@@ 4.1 Baseline methods
We recall that with black-box inference we mean that the target lacks gradients and may be expensive to evaluate and possibly noisy. As few other methods afford post-process and black-box inference of the posterior, we compare vsbq against three baselines: black-box variational inference (bbvi), neural network regression (nnr), and the popular Laplace approximation (laplace). Of these, only nnr is also a post-process method. Further implementation details of vsbq are provided in Supplementary Material C.1.
Black-box variational inference
When the target posterior density is a black-box, the reparameterization trick (Kingma and Welling, 2013) cannot be applied to estimate the elbo gradient for stochastic variational inference. Instead, one needs to resort to other techniques for computing the gradient of the elbo, such as the score function estimator, also known as the REINFORCE estimator (Ranganath et al., 2014), which often has higher variance compared to the reparameterization trick (Gal, 2016; Xu et al., 2019). Specifically, by differentiating Eq. 2, the elbo gradient $\nabla_{\bm{\phi}}\textsc{elbo}\left[q_{\bm{\phi}}\right]$ can be written as,
@@
Neural network regression
For a direct comparison with our method, we develop a post-process inference algorithm based on a deep neural network regression surrogate (nnr) instead of a sparse GP surrogate, otherwise leaving the post-process procedure (Figure 2) as much as possible the same. This is a competitive baseline since deep neural networks exhibit strong regression performance in the presence of a large number of data points (Goodfellow et al., 2016). For the network architecture, we use a multilayer perceptron (MLP) with an input layer of dimension $D$, four hidden layers of $1024$ units, and an output layer for predicting the log-density value. The activation function is chosen to be the rectified linear units (ReLU; Goodfellow et al., 2016, Chapter 6). In addition, we add a negative quadratic mean function to the neural network output to ensure that it represents a valid log-density surrogate function. [^7] An MLP with ReLU activations is a continuous piecewise affine function (Arora et al., 2018), and therefore adding a trainable negative quadratic mean function ensures the integrability of the exponentiated surrogate. The negative quadratic mean function is the same as the one used for the (sparse) GP (see Eq. 17). In total, the surrogate function $g$ is:
@@
$\mathcal{L}(\mathbf{w})=\sum_{n=1}^{N}\frac{\left(g(\mathbf{x}_{n};\mathbf{w})-y_{n}\right)^{2}}{\sigma_{\text{tot}}^{2}(\mathbf{x}_{n},y_{n})},$ (35)
where $\sigma_{\text{tot}}^{2}(\mathbf{x}_{n},y_{n})=\sigma_{\text{obs}}^{2}(\mathbf{x}_{n})+\sigma_{\text{shape}}^{2}(\Delta y_{n})$, as per noise shaping. [^8] We found that noise shaping empirically helped stabilize the neural network training. We optimize the neural network parameters by minimizing the objective in Eq. 35 with the AdamW optimizer (Loshchilov and Hutter, 2019). For regularization, we considered the ‘weight decay’ hyperparameter $\alpha\in\{0,0.01,0.1\}$, selecting the best neural network surrogate based on the loss on a split validation dataset. Finally, we use stochastic variational inference with the reparameterization trick (Kingma and Welling, 2013) to compute the approximate posterior. This part is the same as vsbq except that the expected log joint of the surrogate in Eq. 2 is approximated via Monte Carlo samples rather than calculated exactly via sparse Bayesian quadrature. For more details, see Supplementary Material C.4.
Laplace approximation
@@ 4.2 Procedure and metrics
Procedure
The optimization trace points and corresponding log-density values are collected as the training dataset for vsbq and nnr. For each problem, we repeated the entire optimization procedure – each with multiple MAP estimation runs, as explained above – ten times with different random seeds. This yielded ten different training sets per problem, used to assess the robustness and reliability of the methods. The number of inducing points for the sparse GP is set to $100D$. The number of mixture components $K$ is 50 for both vsbq and nnr. For bbvi, as mentioned in Section 4.1, a budget of $10\times 3000D=30000D$ target density evaluations per random seed is allocated for stochastic optimization. In the case of a noisy target, we further vary the observation noise level to study the noise sensitivity of the methods. For the Laplace approximation, we first find the MAP point by transforming the parameter space to unbounded if needed, via a nonlinear mapping. [^9] For the Laplace approximation to be valid, it is necessary to find the MAP estimate in the unbounded space subsequently used to compute the multivariate normal approximation; this is particularly important if the mode is close to the bounds. We subsequently compute the Hessian matrix via adaptive numerical differentiation (Brodtkorb and D’Errico, 2022). Finally, we compute the performance metrics as detailed below.
Metrics
@@
    where $D_{\text{KL}}\left(p||q\right)$ is the Kullback-Leibler divergence between distributions $p$ and $q$ and $\mathcal{N}[p]$ is a multivariate normal distribution with mean equal to the mean of $p$ and covariance matrix equal to the covariance of $p$ (and same for $q$). [^10] In contrast to the definition in Acerbi (2020), we normalize the GsKL metric by the number of dimensions $D$. Eq. 37 can be expressed in closed form in terms of the means and covariance matrices of $p$ and $q$. For reference, two Gaussians with unit variance and whose means differ by $\sqrt{2}$ (resp., $\frac{1}{2}$) have a GsKL of 1 (resp., $\frac{1}{8}$). As a rule of thumb, we consider a desirable target to have GsKL less than $\frac{1}{8}$.
@@ 4.3 Synthetic problems
Two Moons bimodal posterior
@@
Multivariate Rosenbrock-Gaussian
@@ 4.4 Real-world models
Noisy likelihood evaluations
@@
Bayesian timing model
@@
Figure 5: Bayesian timing model. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML $=1$, MMTV $=0.2$, GsKL $=1/8$). vsbq performs well across all noise levels and can outperform (noiseless) laplace even under high log-likelihood noise, whereas nnr demonstrates less robustness with several failed runs. bbvi methods exhibit similar performance to each other and are above the metric thresholds.
@@
Multisensory causal inference
@@
Figure 6: Multisensory causal inference model. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the multisensory causal inference model. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML $=1$, MMTV $=0.2$, GsKL $=1/8$). vsbq performs well across low to moderate noise levels and outperforms (noiseless) laplace even at high noise levels. nnr also demonstrates strong performance across all noise levels. bbvi methods perform similarly to each other and are above the thresholds, except for the MoG($K=50$) at low noise levels.
@@ 4.5 Additional analyses
We summarize here the results of a number of additional analyses, with further details reported in the Supplementary Material.
MAP estimation with Bayesian Adaptive Direct Search (BADS)
@@
(a) Bayesian timing model ($\sigma_{\text{obs}}=3$).
(b) Multisensory causal inference model ($\sigma_{\text{obs}}=0$).
Figure 7: Sensitivity to the number of target evaluations. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the number of target evaluations, for two benchmark problems. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML $=1$, MMTV $=0.2$, GsKL $=1/8$). Across both tasks, vsbq demonstrates better robustness to the number of evaluations compared to nnr.
Sensitivity to the number of target evaluations
The number of target density evaluations is a critical factor in determining coverage over the posterior distribution, and therefore, would naturally impact the performance of both nnr and vsbq. Throughout the experiments, we used a fixed number of target density evaluations, $3000D$, which corresponds to at least two complete MAP optimization trajectories for all benchmark problems. To investigate the sensitivity of our method to the number of target density evaluations $N$, we further conducted experiments across three evaluation budgets: $N\in\{1000D,2000D,3000D\}$, where $D$ is the dimensionality of the parameter space.
Figure 7 summarizes the results on two representative benchmarks: the noisy Bayesian timing model ($\sigma_{\text{obs}}=3$) and the multisensory causal inference model ($\sigma_{\text{obs}}=0$). We find that vsbq is robust to changes in the number of evaluations, whereas nnr is more sensitive and tends to perform significantly worse with fewer evaluations. Additional results for the bads optimizer are reported in Supplementary Material Section C.5. In contrast to cma-es, bads can lead to poorer coverage of the posterior, as discussed in Supplementary Section C.5, making vsbq more sensitive to the number of evaluations in this case. Nonetheless, vsbq consistently outperforms nnr across most tested settings.
@@
In the earlier paragraphs, we presented results obtained by approximating the target posterior density from the surrogate of the log density (Figure 1) via variational inference. A natural alternative is to instead run mcmc on the surrogate log-density (Rasmussen, 2003; Nemeth and Sherlock, 2018; Järvenpää et al., 2021; El Gammal et al., 2023). In Supplementary Material C.8, we provide experimental results showing that mcmc can underperform in this setting and discuss the reasons.
@@ 5.2 Limitations and future work
Arguably, post-process inference via vsbq is constrained to low-dimensional problems (e.g., up to ten parameters) due to several reasons. As the dimension increases, pre-existing evaluations from MAP optimizations become less likely to cover the majority of the posterior mass (Vershynin, 2018, Chapter 3.3.3), while sufficient coverage of the posterior is essential for vsbq to perform well. Moreover, the MAP estimate may become not meaningful or even ill-defined in certain settings, e.g., when the likelihood is unbounded (Gelman et al., 2013). While a larger number of evaluations might expand the range of tractable dimensions, the curse of dimensionality poses challenges in approximating a function without additional structural assumptions. A potential solution to address the issue of lack of coverage is active sampling, that is, the ability to acquire new log-density evaluations where needed (Bliznyuk et al., 2008; Osborne et al., 2012; Acerbi, 2020; Järvenpää et al., 2021; De Souza et al., 2022). Active sampling would need one or multiple rounds of ad-hoc function evaluations and subsequent surrogate model updates, which would substantially increase the post-processing time (De Souza et al., 2022). Finding efficient methods for active learning for post-process inference is left for future research. Despite the above limitations, the relatively cheap cost of our method makes it still valuable for quickly constructing a tractable initial approximation of the posterior, potentially useful to inform subsequent runs of mcmc or other inference methods (Zhang et al., 2022).
@@
An important feature of approximate inference methods is the availability of diagnostics to assess the reliability of inference (Vehtari et al., 2021; Yao et al., 2018). A general-purpose inference diagnostic consists of posterior-predictive checks, i.e., testing that synthetic data generated from parameters sampled from the posterior are compatible with the actual data (Gelman et al., 2013). As an additional diagnostic, vsbq provides the standard deviation of the elbo, $\textsc{elbo}_{\text{sd}}$, which can be calculated via sparse Bayesian quadrature (Eq. 23). $\textsc{elbo}_{\text{sd}}$ reflects the uncertainty of the sparse GP prediction in regions where the variational posterior has non-negligible mass. Therefore, it can serve as a useful diagnostic in that solutions with $\textsc{elbo}_{\text{sd}}\gg 1$ should not be trusted. However, as is often the case with inference diagnostics, a small $\textsc{elbo}_{\text{sd}}$ does not guarantee the validity of the approximate variational posterior (Acerbi, 2018). As a practical validation approach, we recommend running vsbq multiple times by dropping for example 20% of the training set and checking the consistency of the approximate posteriors via a form of cross-validation. Visualization of the posterior together with the locations of the evaluated points is also helpful for validating that the approximate posterior is supported by actual evaluations, as opposed to escaping from the training points region, leading to ‘hallucinated’ posterior regions (De Souza et al., 2022). We further discuss this ‘hallucination’ problem in Supplementary Material C.8. Finally, if additional exact log-density evaluations are possible, one can leverage Pareto smoothed importance sampling (Vehtari et al., 2024) for correcting and validating the approximate posterior (Yao et al., 2018).